OpenStudy (anonymous):

find all the zeros of f(x)=x^3-3x^2+4x-12

OpenStudy (tamtoan):

set f(x) = 0, and what factor of x^3 - 3x^2 ? what factor of 4x - 12 ? can you factor those ?

OpenStudy (anonymous):

@tamtoan the zero is -12 are there any other zeros that I don't see. What do I do with the factor of (x^3-3x^2) and (4x-12).

OpenStudy (tamtoan):

factor them, and then group them, and set those 2 factors equal zero, then you will find the solution...can you tell me the factors?

OpenStudy (tamtoan):

find all zeros in a function means ..that you have to set that function equal to zero ...ie...find all value of x such that f(x) = 0

OpenStudy (anonymous):

the factors are x^2(x-3) and 4(x-3) so I set them both to 0 and find the solution

OpenStudy (tamtoan):

ok now , x^3-3x^2+4x-12 = 0 so you have x^2(x-3) + 4(x-3) = 0 can you do one more step factoring from here and let me know?

OpenStudy (anonymous):

no

OpenStudy (tamtoan):

look at both of the terms, what do they have in common?

OpenStudy (anonymous):

x-3

OpenStudy (tamtoan):

ok if x-3 is the common term, then you can factor that out, correct? ...can you show me how you factor that?

OpenStudy (anonymous):

nope i dont know how to factor out

OpenStudy (tamtoan):

ok...let me do this for you, if anything you don't understand, ask a question again ok? find all the zeros of f(x)=x^3-3x^2+4x-12 set f(x) = 0, then x^3-3x^2+4x-12 = 0 x^2(x-3) + 4(x-3) = 0 since both terms have x-3, factor that out, you have: (x-3)(x^2 + 4) = 0 now 2 factors multiple each other equal to zero means either one of them is zero. so you will have : x - 3 = 0 or, x^2 + 4 = 0 the first one, you can solve easy right? the second one, if you already learn complex number, you can solve that too, let me know your answer ?

OpenStudy (anonymous):

thanks

OpenStudy (tamtoan):

you got it right? , can you please tell me what are the zeros ? ie. value of x ?